OpenStudy (anonymous):

what is the LCD of 3 7 9 2

OpenStudy (zzr0ck3r):

least common denominator?

OpenStudy (zzr0ck3r):

should that be \[\frac{3}{7}and\frac{9}{2}?\]

OpenStudy (zzr0ck3r):

@Lady_Bella

OpenStudy (mathstudent55):

Find the prime factors of each number: 3: 3 7: 7 9: 3 x 3 = 3^2 2: 2 For the least common denominator, use common and not common factors with highest exponent, so LCD = 7 x 3^2 x 2 = ?

OpenStudy (zzr0ck3r):

im confused @mathstudent55 the least common denominator or 3,7,9,2 is 1 the least common denominator of (3/7) and (9/2) is 14 what are you talking about that equals 126?